# MCCoder: Streamlining Motion Control with LLM-Assisted Code Generation and Rigorous Verification

### v1

@@ Abstract
Large Language Models (LLMs) have shown considerable promise in code generation. However, the automation sector, especially in motion control, continues to rely heavily on manual programming due to the complexity of tasks and critical safety considerations. In this domain, incorrect code execution can pose risks to both machinery and personnel, necessitating specialized expertise. To address these challenges, we introduce MCCoder, an LLM-powered system designed to generate code that addresses complex motion control tasks, with integrated soft-motion data verification. MCCoder enhances code generation through multitask decomposition, hybrid retrieval-augmented generation (RAG), and self-correction with a private motion library. Moreover, it supports data verification by logging detailed trajectory data and providing simulations and plots, allowing users to assess the accuracy of the generated code and bolstering confidence in LLM-based programming. To ensure robust validation, we propose MCEVAL, an evaluation dataset with metrics tailored to motion control tasks of varying difficulties. Experiments indicate that MCCoder improves performance by 11.61% overall and by 66.12% on complex tasks in MCEVAL dataset compared with base models with naive RAG. This system and dataset aim to facilitate the application of code generation in automation settings with strict safety requirements. MCCoder is publicly available at https://github.com/MCCodeAI/MCCoder.
@@ I INTRODUCTION
Figure 1: Naive RAG code generation vs MCCoder. Code generation with Naive RAG has a low pass rate, requiring extensive debugging and posing safety concerns. In contrast, MCCoder generates code through task decomposition, correction, and simulation with soft-motion, ensuring data verification and safe machine deployment.
Motion control, a cornerstone of automation, has significantly advanced industrial processes since the advent of NC machines in the 1950s. By precisely controlling actuators, it has revolutionized manufacturing, from early CNC machines to today’s sophisticated AI-driven robots and semiconductor equipment. While motion control encompasses various programming methods (e.g., CAD/CAM, PLC, robot teaching), this paper focuses on the complex domain of motion API invocation. Today, software engineers in this industry still manually program due to its complexity and safety concerns. Some private motion libraries have over 1,000 APIs with various arguments, creating a steep learning curve and making reliance on automatic programming tools difficult. Additionally, engineers must code and iteratively test with electrical and mechanical components to ensure optimal motion performance and eliminate potential hazards for humans and machines during execution.
The recent spotlight on LLM-based code generation has unveiled impressive ability to address programming challenges [1]. Although traditional code generation in automation control often relied on model-driven or rule-based approaches, recent advancements have started to explore LLM-based methods. For example, a retrieval-augmented approach [2] was proposed to generate IEC 61131-3 programs, while LLM4PLC [3] used user feedback and external verification tools to guide the generation of LLM-based PLC codes.
@@
To address these limitations, we propose the MCCoder, which utilizes LLMs to generate Python code for motion control based on natural language instructions. This system employs a private motion library to decompose complex tasks, generate code, and use soft-motion for simulation running and data logging for verification. In addition, we build the MCEVAL dataset to evaluate the generated code. Figure 1 highlights a comparation between LLM-based code generation with naive RAG and LLM-powered MCCoder. Experiments demonstrated that MCCoder enhances code generation performance, achieving an overall improvement of 11.61% and a significant increase of 66.12% in complex tasks in the MCEVAL dataset.
@@
  • Code Generation System: We introduce MCCoder system that effectively generates Python code for motion control using LLMs. It incorporates task decomposition, retrieval, code generation, soft-motion, self-correction and data verification to ensure code safety and effectiveness in addressing complex real-world programming challenges. Soft-motion is carefully selected as a flexible motion controller integrated with LLMs for rapid simulation, execution, self-correction, and rigorous verification.
  • Evaluation Dataset: Current code generation datasets and metrics are not designed for the motion control domain. We present MCEVAL, the first benchmark dataset and metrics specifically tailored for evaluating LLM-powered motion control code. This resource fills a gap in the field.
  • Benchmarking: Through extensive experiments comparing state-of-the-art LLMs with MCCoder, we offer insights into their performance and evaluate their suitability for deployment in industrial motion control applications.
@@ II PRELIMINARY
MCCoder utilizes a soft-motion architecture to generate Python code for motion control by invoking APIs from a private library.
@@ II-A Soft-motion
The soft-motion controller has a flexible and lightweight architecture that runs on a general PC without the need for specialized hardware. It uses a dedicated CPU core with a real-time OS and software motion algorithms to deliver performance comparable to high-end hardware controllers. MCCoder sends control code to the soft-motion engine through a DLL. Soft-motion executes the code in a simulation engine first, allowing for data logging and verification before deployment. Additionally, its PC-based software architecture makes soft-motion ideal for integrating new functions, including LLMs and other AI technologies.
II-B Control Code and Private Library
Unlike standardized PLC programming like IEC 61131-3, control codes programming in Python are more flexible and difficult to apply a fixed template. In this research, we use a private motion library in the soft-motion controller with more than 1,000 APIs capable of controlling up to 128 axes and 256 independent task channels. These APIs encompass most motion control functionalities, such as fieldbus communication, digital and analog IO control, and various trajectory and position/velocity/current control for servo axes. The private library also includes hundreds of sample codes. Section II-B shows a simple code example to move Axes to positions, including the INITIALIZE, main control code and CLOSE procedure.
Example 1. A Motion Control Code Demo   Inputs:   $Axes$ = [$A_{i}$] // Axes objects from LLMs   $IO_{Ins}$ = [$I_{i}$]; // Input objects from LLMs   $IO_{Outs}$ = [$O_{i}$] // Output objects from LLMs   procedure INITIALIZE():       $MCLib.CreateDevice()$       $MCLib.StartCommunication()$       $MCLib.SetandStartLog(Axes,IO_{Ins},IO_{Outs})$   // Main control code   for $i$ in $Axes$ do       $posCommand=MotionPosCommand()$      $posCommand.profile.type=ProfileType_{i}$      $posCommand.axis=i$      $posCommand.target=Pos_{i}$      $MCLib.motion.StartPos(posCommand)$      $MCLib.motion.Wait(i)$   end for   procedure CLOSE():           $MCLib.CloseDevice()$   end procedure   Return: Success or error code
@@ III METHODOLOGY
We developed the MCCoder system to tackle the code generation challenges in motion control. Figure 2 shows an overview of the MCCoder system. It involves six modules: task decomposition, retrieval, code generation, soft-motion, self-correction and data verification.
Figure 2: Overview of the MCCoder system. It integrates multitask decomposition, hybrid retrieval, code generation, and self-correction with a private motion library to enhance performance, offering motion data visualization for execution verification before deployment. Soft-motion is a key component in the system to streamline motion control with LLM-assisted code generation with rapid execution and self-correction cycle as a rigorous verification, which is critical for control code programming.
@@ III-A Task Decomposition
The MCCoder system receives user questions as a motion task. It could be a single task or a multitask composing with many sub-tasks. This module will read the task and decompose them to be a list of sub-tasks.
III-B Retrieval
The retrieval module extracts APIs and expert-created sample codes from the private library, each accompanied by clear instructions. Retrieval methods are commonly employed in code generation to provide relevant samples and contexts. Various Retrieval-Augmented Generation (RAG) techniques are available, including native RAG, advanced RAG, and modular RAG [4]. Given that the library contains uniformly formatted sample codes and APIs, we employ Sparse Retrieval (BM25), Dense Retrieval (embedding-based VectorStores), and a re-ranker to match user queries with the most relevant content, setting the top-$k$ as 10.
III-C Code Generation
The code generation model processes sub-tasks and retrieved sample codes to generate sub-codes based on prompts, repeating this for m tasks. It then combines the user question and sub-codes to generate the final code. A critical but often overlooked aspect is prompt engineering. To consistently refine the prompts, we iterated and established 8 guidelines to help models avoid common mistakes.
@@ III-D Soft-Motion
Soft-motion includes a real-time engine for operating actual machines and a simulation engine. MCCoder sends the final code to the soft-motion controller, which runs the code in the simulation engine and returns execution feedback.
@@ III-E Self-Correction
If the soft-motion engine returns a syntax or API error, MCCoder uses the error to retrieve relevant documents, regenerate code and run it again for self-correction iteratively. If no error is detected, soft-motion logs the entire trajectory data for subsequent verification.
@@ III-F Data Verification
We can design a 3D model of a machine and import the model file into the 3D simulator, which displays the motion in real-time, providing users with an intuitive verification when it runs. Additionally, logging data is recorded every 1ms by the soft-motion engine as the process data. These features are essential to build user trust in an AI-driven system through visual and data-driven verification.
@@ IV-A Construction
MCEVAL comprises 116 tasks that were chosen to provide a diverse but manageable set of motion control programming challenges. It offers substantial diversity, covering most common used motion control functions in the library. These functions include point-to-point motion, linear, circular, and helical interpolation, splines, short-segment look-ahead, compensation, and event-driven interactions of axes and IOs under various profiles, as well as their combinations. Each task includes a TaskId, Instruction, CanonicalCode, and Difficulty. The properties of these tasks are detailed in Table I.
To mimic a human engineer’s problem-solving capabilities, we structured the tasks into three difficulty levels. These levels are differentiated by sample code, primarily demonstrating single API function calls and not covering all argument combinations (e.g., there are 19 acceleration profiles). The examples are shown in Sections IV-A and IV-A. Difficulty Level 1 involves invoking similar APIs as in the sample codes but with different numerical arguments, which should be easy for most LLMs if the retrieved sample codes are correct. Difficulty Level 2 uses arguments not shown in the sample codes, making it more challenging. LLMs might make mistakes but should utilize error information or retrieved content for self-correction. Difficulty Level 3 involves combining up to 8 different motion tasks. Consolidating more than 8 tasks into a single instruction may exceed input or output token limits for most models. These levels emulate real-world scenarios from easy to complex motion tasks.
@@
Difficulty Level 1 61 267 2268
Difficulty Level 2 32 289 1351
Difficulty Level 3 23 406 2926
Total 116 273 2145
Example 2. Sample Code   Instruction: Move Axis 2 to 100, velocity as 1000.   Return:       $posCmd=MotionPosCommand()$       $posCmd.profile.type=Trapezoidal$       $posCmd.axis=2$       $posCmd.target=100$       $posCmd.profile.velocity=1000$       $MCLib.motion.StartPos(posCmd)$
[…]
@@ V-A Experimental Settings
We access closed-source models through APIs and open-source one on a local server with eight NVIDIA GeForce RTX 3090 GPUs and Intel(R) Xeon(R) Gold 5218R CPU @ 2.10GHz. A soft-motion controller and 3D simulator were installed on a Windows 11 PC, ready for code execution and simulation.
@@ V-B Base Models
LLMs evolve rapidly, continuously setting new state-of-the-art benchmarks. For this study, we selected the latest closed-source and open-source LLMs as base models. Given the average token length of 1,424 in the sample codes, we opted for models with relatively large context windows and output lengths, as detailed in Table II. GPT-4o, GPT-4o-mini, Lama-3.1-70b, and DeepSeek-Coder-v2 were selected for their proven performance in code generation benchmarks.
@@
Model Context Window Max Output Tokens Active Params
gpt-4o 128,000 4,096 -
gpt-4o-mini 128,000 16,384 -
llama-3.1-70b 128,000 2,048 70B
deepseek-coder-v2 128,000 4,096 21B
@@ V-C Baseline
Since the private library is not in pre-trained LLM data, zero-shot evaluation is inapplicable. We established a baseline using the base model with naive RAG and top-$k$ = 10, comparing it to the base models integrated into the MCCoder system.
@@ V-D Experimental Results
TABLE III: Experiments results of various models in MCEVAL datasets. For each model, the baseline naive RAG and the MCCoder system were compared. Metrics include FTPR (MatchEndPoints) and FTPR (DTW), with the latter being more stringent. The overall first-time pass rate was assessed and performance was evaluated at three difficulty levels. L1, L2, and L3. The best performance are boldfaced for Naive RAG, and underlined for MCCoder.
@@
gpt-4o Naive RAG 67.24 78.69 59.38 47.83 53.45 63.93 46.88 34.78
MCCoder 75.86 83.61 65.62 69.57 67.24 73.77 59.38 60.87
gpt-4o-mini Naive RAG 60.34 75.41 46.88 39.13 43.97 52.46 37.50 30.43
MCCoder 66.38 75.41 53.12 60.87 49.14 55.74 37.50 47.83
llama-3.1-70b Naive RAG 59.48 73.77 50.00 34.78 48.28 59.02 40.62 30.43
MCCoder 64.66 68.85 59.38 60.87 50.00 52.46 50.00 43.48
deepseek-coder-v2 Naive RAG 67.24 77.05 62.50 47.83 56.90 68.85 50.00 34.78
MCCoder 75.86 80.33 65.62 78.26 61.21 67.21 53.12 56.52
We compared naive RAG and MCCoder using FTPR (MatchEndPoints) and FTPR (DTW) metrics. The overall pass rates and three difficulty levels were also evaluated for each model, as shown in Table III. We will present results with research questions.
V-D1 RQ1. Does MCCoder imporve the code generation performance?
MCCoder yielded an overall FTPR improvement of 11.61%. While this might appear modest at first glance, a deeper analysis reveals a more substantial impact at higher difficulty levels. At Level 3, the improvement increases to 66. 12%, demonstrating MCCoder’s exceptional ability to tackle complex real-world challenges.
@@
GPT-4o demonstrated the best overall performance, particularly with MCCoder. However, DeepSeek-Coder-v2 exhibited very strong performance comparable to GPT-4o and even outperformed it in naive RAG, which is surprising given its 24 billion active parameters. In the HumanEval evaluation, DeepSeek-Coder-v2 scored 90.2, close to GPT-4o’s 91.0 [8]. Our experiments show similar results. GPT-4o-mini and Lama-3.1-70B performed slightly weaker.
@@
In general-purpose Python programming, verification typically focuses on execution results via unit tests. In motion control, however, engineers are concerned with the entire execution process, including the trajectory and endpoints of the objects under control to confirm if it is correct and safe. LLMs’ inferencing and code generation capabilities are often perceived as a black-box, raising transparency concerns. Therefore, traceable data that clearly represent the entire process is essential. For each task execution, MCCoder generates a logging file, capturing the entire process as time-series data for evaluation. In real-world scenarios, 1D-3D plots of Axes and IOs, along with real-time 3D simulator of the machine model, provide a multimodal, traceable method for verifying whether the trajectory and endpoints align with expectations. MCCoder provides traceability that is crucial for building trust and enabling deployment, as presented in Figure 3.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated significant potential in code generation. However, in the factory automation sector—particularly motion control—manual programming, alongside inefficient and unsafe debugging practices, remains prevalent. This stems from the complex interplay of mechanical and electrical systems and stringent safety requirements. Moreover, most current AI-assisted motion control programming efforts focus on PLCs, with little attention given to high-level languages and function libraries. To address these challenges, we introduce MCCoder, an LLM-powered system tailored for generating motion control code, integrated with a soft-motion controller. MCCoder improves code generation through a structured workflow that combines multitask decomposition, hybrid retrieval-augmented generation (RAG), and iterative self-correction, utilizing a well-established motion library. Additionally, it integrates a 3D simulator for intuitive motion validation and logs of full motion trajectories for data verification, significantly enhancing accuracy and safety. In the absence of benchmark datasets and metrics tailored for evaluating motion control code generation, we propose MCEVAL, a dataset spanning motion tasks of varying complexity. Experiments show that MCCoder outperforms baseline models using Advanced RAG, achieving an overall performance gain of 33.09% and a 131.77% improvement on complex tasks in the MCEVAL dataset. MCCoder is publicly available at https://github.com/MCCodeAI/MCCoder.
@@ I INTRODUCTION
Figure 1: Comparison between Advanced RAG and LLM-assisted MCCoder. Code generation with Advanced RAG has a low pass rate and efficiency, requiring extensive debugging and posing safety concerns. In contrast, MCCoder generates code through task decomposition, self-correction, and soft-motion with simulator, ensuring data verification and safe machine deployment.
Motion control, a fundamental element of factory automation, has greatly improved industrial processes, evolving from the introduction of numerical control (NC) machines in the 1950s to the advanced AI-powered robots and semiconductor equipment of today.While motion control spans a range of programming approaches (e.g., CAD/CAM, PLC, and robot teaching), this paper concentrates specifically on the domain of motion API invocation. Presently, software engineers in this field primarily depend on manual programming and debugging of motion control systems, driven by the intricate nature of manufacturing processes and critical machinery safety considerations. Certain motion libraries include over 1,000 APIs with override arguments and more than 100 synchronous motion channels, creating a steep learning curve and limiting the practicality of automated programming tools. In addition, engineers must iteratively debug in coordination with electrical and mechanical components to enhance motion performance and address potential hazards to both human operators and machinery during operations.
The recent spotlight on LLM-based code generation has unveiled impressive ability to address programming challenges [1]. Although traditional code generation in automation control often relied on model-driven or rule-based approaches, recent advancements have begun to explore LLM-based methods. For example, a retrieval-augmented approach [2] was proposed to generate IEC 61131-3 programs, while LLM4PLC [3] used user feedback and external verification tools to guide the generation of LLM-based PLC codes.
@@
To address these limitations, we propose MCCoder, which utilizes LLMs to generate Python code for motion control based on natural language instructions. This system employs a motion library to decompose complex tasks, generate code, and use soft-motion for simulation running and data logging for verification. In addition, we build the MCEVAL dataset to evaluate the generated code. Figure 1 highlights a comparison between code generation with Advanced RAG and LLM-assisted MCCoder. Experiments demonstrated that MCCoder improves code generation performance, achieving an overall improvement of 33.09% and a significant increase of 131.77% in complex tasks in the MCEVAL dataset.
@@
  • MCCoder System: MCCoder is an AI-assisted code generation and verification system specifically designed for motion control in factory automation. It integrates task decomposition, hybrid retrieval, code generation, soft-motion, self-correction, and data verification to ensure code safety and effectiveness in handling complex real-world programming challenges. Soft-motion with LLM plays a central role, transforming the traditional approach of manual programming or LLM-based coding, and human debugging into an automated workflow.
  • Evaluation Dataset and Metrics: Existing code generation datasets and metrics are not designed for the motion control domain. We introduce MCEVAL, the first benchmark dataset and evaluation metrics specifically tailored for LLM-assisted motion control code. MCEVAL comprises 186 motion control programming tasks, ranging from simple to complex, covering most common functionalities and domains in motion control. It carefully considers function diversity, potential ambiguities, and interactions among multiple subtasks, while its evaluation metrics account for both motion endpoints and trajectory assessment. MCEVAL fills the gap in general-purpose code generation benchmarks and the motion control field, serving as a valuable reference for professionals.
@@ II PRELIMINARY
MCCoder utilizes a soft-motion architecture to generate Python code for motion control by calling APIs from WMX3 motion library.
@@ II-A Soft-motion
A motion controller is typically used to control the motion of servo motors, axes, and I/Os in machines, relying on dedicated hardware architecture. In contrast, a soft-motion system features a flexible and software-based architecture that runs on a general-purpose PC without requiring specialized hardware. It leverages a CPU core with a real-time OS and complex motion algorithms, ensuring high performance and scalability.
MCCoder interacts with the soft-motion via a DLL. The soft-motion system first executes the control code in a simulation engine for data logging and verification. Once validated, the real-time engine transmits control commands to the machine through cyclic fieldbus communication. Moreover, its PC-based software architecture makes soft-motion particularly well suited for integrating advanced functionalities, such as LLMs and other AI technologies.
II-B Control Code and Motion Library
Unlike standardized PLC programming, such as IEC 61131-3, control code programming in Python offers greater flexibility but is harder to constrain within a fixed template. In this research, we utilize the well-known WMX3 motion library within a soft-motion controller, which provides over 1,000 APIs capable of managing up to 128 axes and 256 independent task channels. These APIs cover a wide range of motion control functionalities, including fieldbus communication, digital and analog I/O control, and diverse trajectory, position, velocity, and current control for servo axes. Notably, these API functions exhibit a complex structure: the same function name may correspond to multiple overrides and parameter classes, requiring even experts with extensive industry experience to carefully distinguish their differences, combinations, and application scenarios—a significant challenge for AI-driven programming. Section II-B presents a sample code demonstrating how to move axes to specified positions, encompassing the INITIALIZE, MOTION, and CLOSE procedures.
Example 1. Motion Control Sample Code   $Axes$ = [$A_{i}$]   $IO_{Ins}$ = [$I_{i}$]   $IO_{Outs}$ = [$O_{i}$]   INITIALIZE():   $MCLib.CreateDevice()$   $MCLib.StartCommunication()$   $MCLib.SetandStartLog(Axes,IO_{Ins},IO_{Outs})$   MOTION():   for $i$ in $Axes$ do      $posCommand=MotionPosCommand()$      $posCommand.profile.type=ProfileType_{i}$      $posCommand.axis=i$      $posCommand.target=Pos_{i}$      $MCLib.motion.StartPos(posCommand)$      $MCLib.motion.Wait(i)$   end for   CLOSE():   $MCLib.CloseDevice()$   RETURN: Success or error code
@@ III METHODOLOGY
We developed the MCCoder system to tackle the code generation challenges in motion control. Figure 2 shows an overview of the MCCoder system. It involves six modules: task decomposition, hybrid retrieval, control code generation, soft-motion, self-correction and data verification.
Figure 2: Overview of the MCCoder system. It integrates task decomposition, hybrid retrieval-augmented code generation, simulation, and self-correction with soft-motion system to enhance control code programming. It employs instinct-driven and data-driven verification before machine execution and deployment, greatly improving the efficiency and safety of control code generation.
@@ III-A Task Decomposition
The Task Decomposition Module, powered by an LLM, accepts user questions phrased as motion control tasks in natural language. These questions may involve multiple interconnected control tasks with complex dependencies and mutual triggering relationships. To enhance the LLM’s comprehension and facilitate the retrieval of pertinent example code and documentation, this module breaks down the question into discrete subtasks for subsequent processing. Additionally, it extracts axis numbers and input/output identifiers from the question, preparing the necessary preprocessing for the downstream soft-motion system.
III-B Hybrid Retrieval
The retrieval module aims to efficiently locate relevant information from extensive documentation and sample code. Thousands of pages of explanatory documents and code examples are chunked separately, using distinct strategies to preserve code integrity. Prior studies have introduced various RAG techniques, such as native RAG, advanced RAG, and modular RAG [4]. In our case, accurately retrieving correct API overrides and example code is challenging due to high similarity among motion APIs. Thus, we employ Sparse Retrieval (BM25), Dense Retrieval (embedding-based VectorStores), and a re-ranker to select the top-$k$ (set as 6) results, effectively matching user queries with the most relevant content.
III-C Control Code Generation
This process employs an LLM combined with retrieved motion-specific sample code and documentation to generate control code. Prompt engineering is essential for enforcing proper code formatting, standards, and critical constraints. The generated code undergoes preprocessing, including initialization steps such as enabling the corresponding servo axes, configuring IO, and setting up logging prior to execution. Termination logging and process closure instructions are appended after the code. This ensures the code is properly formatted and runnable before deployment to the soft-motion system.
@@ III-D Soft-Motion
The soft-motion system comprises a real-time engine for controlling physical machines and a simulation engine. MCCoder transmits the finalized code to the soft-motion system, which executes it in the simulation engine and provides feedback, including correct results or error messages for self-correction. Additionally, the simulation engine logs motion data from axes and IO at 1ms intervals throughout execution, facilitating process verification and endpoint validation.
@@ III-E Self-Correction
If the soft-motion system detects a syntax or API error, MCCoder retrieves relevant documentation, regenerates the code, and re-executes it through an iterative self-correction process. Once execution is error-free, the soft-motion system logs the full trajectory data for subsequent verification. This synergy establishes a closed-loop feedback mechanism, enabling iterative code refinement and secure validation.
@@ III-F Data Verification
The 3D model of a machine can be imported into a 3D simulator, enabling instinct-driven verification by visualizing real-time motion during the execution of motion code in the soft-motion system and alerting users to potential collisions. Additionally, the soft-motion engine performs data-driven verification by logging comprehensive execution data every 1ms, capturing all axis and IO details to facilitate fine-grained analysis and ensure process accuracy.
@@ IV-A Construction
MCEVAL consists of 186 tasks selected to offer a diverse range of motion control programming challenges. It encompasses the most commonly used motion control functions in the library, spanning from simple to complex. These functions cover point-to-point motion, linear, circular, and helical interpolation, splines, short-segment look-ahead, compensation, and event-driven interactions between axes and IOs across various profiles, along with their combinations and dependencies. The tasks also account for inaccuracies in user input. Each task is defined by a TaskId, Instruction, CanonicalCode, and Difficulty, with their properties detailed in Table I.
To replicate a human engineer’s problem-solving skills, we categorized the tasks into three difficulty levels, with examples provided in Section IV-A. Difficulty Level 1 involves calling standard APIs with straightforward arguments as outlined in the documentation. Difficulty Level 2 features APIs with more complex parameter types, including easily confused overrides, as well as potential user errors such as typos or improper parameter instructions. LLMs may err here but should leverage error feedback or retrieved content for self-correction. Difficulty Level 3 requires integrating multiple motion tasks, not merely stacking simple ones, but combining them into a cohesive set linked by events and dependencies. Solving these demands robust retrieval of all subtasks and a comprehensive understanding of the overall task. These levels mirror real-world motion control scenarios, progressing from simple to complex.
@@
Difficulty Level 1 84 267 2268
Difficulty Level 2 56 289 1351
Difficulty Level 3 46 406 2926
Total 186 273 2145
Example 2. MCEVAL Dataset   // Difficulty: 1   Instruction: Write Python code to move axis 1 to position 130.2 at a speed of 1060, and acceleration of 11000.     // Difficulty: 2   Instruction: Write Python code to move Axis 9 to the position 90 at a speed of 1000, jerkAccRatio of 0.5, end Velocity of 0, using a Jerk-Ratio profile.     // Difficulty: 3   Instruction: Write Python code to set the input event to monitor if the DistanceToTarget of Axis 3’s movement is 500, then move Axis 1 to the position -200 at a speed of 1000. Move Axis 3 to 1200…
@@ V-A Experimental Settings
We established a demo machine, as depicted in Figure 3, to evaluate MCCoder’s capabilities. The setup comprises a Windows PC running the MCCoder system, which interacts with various cloud-based LLMs and features a soft-motion system with both simulation and real-time engines, enabling code validation on the machine. The demo setup also integrates precision linear motor stages with five axes and a six-axis collaborative robot (Cobot from Schneider). The electrical hardware includes EtherCAT servo drives for precise motion control and an I/O module featuring 16 digital inputs and 16 digital outputs, all of which facilitate comprehensive validation of the generated code.
Figure 3: MCCoder Demo Machine.
@@ V-B Base Models
LLMs are advancing rapidly, consistently setting new state-of-the-art benchmarks. For this study, we selected the latest chat and reasoning LLMs as base models, as outlined in Table II. GPT-4o, o3-mini, DeepSeek-V3, and DeepSeek-R1 were chosen for their proven excellence in code generation tasks. Reasoning models were included for their robust logical problem-solving and deep contextual understanding, essential for accurate and efficient automated code generation.
@@
Model Context Window Max Output Tokens Params
gpt-4o 200,000 16,384 -
o3-mini 200,000 100,000 -
DeepSeek-V3 64,000 8,192 671B
DeepSeek-R1 64,000 8,192 671B
@@ V-C Baseline
Since motion control engineers nowadays frequently leverage chat-based models such as ChatGPT for code generation assistance, we designed our baseline to reflect this practical scenario. Specifically, we adopted an Advanced RAG approach, integrating hybrid retrieval and reranking to enhance contextual accuracy. This baseline serves as a reference for comparing the performance of MCCoder on the MCEVAL dataset.
@@ V-D Experimental Results
TABLE III: Experiments results of various models in MCEVAL datasets. For each model, the baseline Advanced RAG and the MCCoder system were compared. Metrics include FTPR (MatchEndPoints) and FTPR (DTW), with the latter being more stringent. The overall first-time pass rate was assessed and performance was evaluated at three difficulty levels. L1, L2, and L3. The best performance are boldfaced for Advanced RAG, and underlined for MCCoder.
@@
gpt-4o Advanced RAG 52.69 70.24 46.43 28.26 63.98 84.52 55.36 36.96
MCCoder 82.80 80.95 80.36 89.13 69.35 69.05 71.43 67.39
o3-mini Advanced RAG 55.38 70.24 55.36 28.26 44.09 60.71 42.86 15.22
MCCoder 72.58 80.95 76.79 52.17 60.75 71.43 67.86 32.61
DeepSeek-V3 Advanced RAG 65.05 88.10 57.14 32.61 53.23 73.81 46.43 23.91
MCCoder 81.72 89.29 75.00 76.09 65.59 72.62 64.29 54.35
DeepSeek-R1 Advanced RAG 66.67 86.90 62.50 34.78 56.45 76.19 51.79 26.09
MCCoder 79.03 90.48 71.43 67.39 66.13 78.57 58.93 52.17
We compared Advanced RAG and MCCoder using FTPR (MatchEndPoints) and FTPR (DTW) metrics. The overall pass rates and three difficulty levels were also evaluated for each model, as shown in Table III. We will present results with three research questions.
V-D1 RQ1. Does MCCoder improve the code generation performance?
Taking FTPR (MatchEndPoints) as an example, MCCoder achieved an overall FTPR improvement of 33.09% compared to the baseline. Specifically, the improvements for Difficulty levels 1, 2, and 3 were 8.99%, 39.33%, and 131.77%, respectively. This indicates that MCCoder significantly outperforms Advanced RAG, especially for more complex motion tasks. In other words, MCCoder demonstrates superior performance compared to the approach where control engineers manually provide sample codes to large models for code generation.
@@
MCCoder achieved the best overall performance when using gpt-4o as the base model, excelling in both matching end points and entire trajectory. The overall ranking of MCCoder’s performance across different base models is: gpt-4o>DeepSeek-V3>DeepSeek-R1>o3-mini. DeepSeek-V3 and R1 delivered results very close to gpt-4o, even surpassing it in some specific sections. However, when self-correction was required due to errors in generated answers, the Chain-of-Thought (CoT) revealed that the reasoning model tended to overthink in these specific motion control tasks. This led to an increase in both redundancy and error rates in the corrected outputs.
Meanwhile, o3-mini demonstrated the weakest overall performance, possibly due to its smaller model size, which may have limited its ability to handle the complexity of control tasks effectively.
@@
In general-purpose Python programming, verification typically focuses on execution results via unit tests. In motion control, however, engineers must evaluate the entire execution process, including trajectory and endpoints, to ensure correctness and safety. Traceable data that clearly represent the entire process is essential.
MCCoder employs two verification approaches: instinct-driven verification through a real-time 3D simulator of the machine model and data-driven verification via logging files that capture the entire execution as time-series data, including both endpoints and trajectories. Additionally, 1D-3D trajectory plots of Axes and IOs provide a multimodal, traceable method to ensure the process aligns with expectations, as shown in Figure 4.